24,32,91,337,1473,?,45829

Answer by Edwin McCravy(20054)

24×1+2³ = 24+8 = 32
32×2+3³ = 64+27 = 91
91×3+4³ = 273+64 = 337
337×4+5³ = 1348+125 = 1473
1473×5+6³ = 7365+216 = 7581  <--missing term
7581×6+7³ = 45486+343 = 45829

Edwin